OpenStudy (anonymous):

Simplify sin3xcos2x-cos3xsin2x

OpenStudy (anonymous):

Can you please help? @myininaya

OpenStudy (anonymous):

looks like the subtraction angle formula for sine

OpenStudy (anonymous):

\[\sin(\alpha -\beta)=\sin(\alpha)\cos(\beta)-\cos(\alpha)\sin(\beta)\] put \(\alpha =3x, \beta=2x\)

OpenStudy (anonymous):

Do is it sinx?

OpenStudy (anonymous):

you get \[\sin(3x-2x)\] or just \[\sin(x)\] that is what i get, yes

OpenStudy (anonymous):

thank you!